# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v1

@@ 1 Introduction
$K_{h}$-minor-free $\operatorname{\mathcal{O}}(h\sqrt{n})$ $\operatorname{\mathcal{O}}(n^{1+\varepsilon})$ [^2] Assumes $h$ is a constant. [KR10]
@@ 2.1 Bounds on $\lambda_{2}(G)$ for Special Graph Classes
For a maximal genus-$g$ graph $G$ with $n$ vertices and maximum degree $\Delta$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\text{poly}(\Delta)\cdot g}{n}$, where $\text{poly}(\Delta)$ denotes a polynomial in $\Delta$.
@@ 3.2 Bounded-Genus Graphs
Kelner’s bound on $\lambda_{2}$ (stated in Theorem 2.3) for maximal bounded-genus graphs also makes use of circle packings [Kel04, Kel06b, Kel06a], and, as we will see, can also be generalized to a bound on $\lambda_{2}^{*}$ and $\gamma^{(1)}$. However, the resulting bound also has a (large) polynomial dependence on the maximum degree $\Delta$. Rather than directly applying this bound, we will apply a new transformation that allows us to almost fully tighten this dependence. Specifically, our new transformation will take a genus-$g$ graph with $n$ vertices and maximum degree $\Delta$, and produce a graph with $O(\Delta\cdot n)$ vertices and constant maximum degree. We will then relate the values of $\gamma^{(1)}$ for the two graphs. In the case of $\lambda_{2}^{*}$ and $\gamma^{(1)}$, this relationship will result in a greatly improved $O(\log\Delta)$ factor for the upper bound.
Also, it is notable that Kelner’s bound on $\lambda_{2}$ does not apply to non-maximal bounded-genus graphs. This does not seem to be due to any significant obstruction, but rather simply because the machinery has not been built for such a bound. We will amend this while tightening and generalizing the bound to $\lambda_{2}^{*}$ and $\gamma^{(1)}$. Although any non-maximal bounded-genus graph can be made maximal simply by adding edges, a naive process may increase some degrees to $\Omega(n)$ for a graph of $n$ vertices. It seems plausible that edges could be added in such a way that the maximum degree remains asymptotically identical, which has been accomplished in the case of planar graphs [KB92, KB97]. However, we will bypass this restrictive construction method entirely with the essentially same method we use to reduce the degree.
@@
Recall that a graph $G$ has genus $g$ if it has a combinatorial embedding (represented as a rotation system) onto a surface of genus $g$. That is, each vertex has a circular ordering on its vertices, forming a locally-planar embedding. A maximal genus-$g$ graph $G$ is also a triangulation. That is, each face defined by the rotation system is a triangle. Importantly for our purposes, this will allow us to make certain local transformations to a graph without changing its genus, so long as the rotation system is respected in these transformations. These transformations will make use of the following structure:
@@
We start with a trivial case whose elimination we will use later: If $\gamma^{(1)}(H)\cdot r\cdot L\gtrsim 1$, then the result holds since $\gamma^{(1)}(G)\lesssim 1$.
@@
We will be able to leverage this lemma in order to turn a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for bounded-genus graphs of constant maximum degree into a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for bounded-genus graphs of arbitrary maximum degree at only a $\log\Delta$ loss. In particular, we will obtain a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for maximal bounded-genus graphs of constant degree by adapting Kelner’s techniques [Kel04, Kel06b, Kel06a] for bounding $\lambda_{2}$. Before exploring these techniques, we first show how to reduce to the case of maximal bounded-genus graphs, which will use a similar technique:
@@
Let $G$ be a graph of genus $g$ with $n$ vertices and maximum degree $\Delta$. Then there is a triangulated (maximal) genus-$g$ graph $H$ with $(\Delta+1)\cdot n$ vertices and maximum degree $O(\Delta)$, so that $\gamma^{(1)}(G)\lesssim(\Delta+1)\cdot\gamma^{(1)}(H)$.
@@
Let $H_{1}$ be the graph that adds $\Delta$ vertices around each vertex to $G$, and ensure that between any two edges in the rotation system around a vertex in $G$, at least one such edge is added. Observe that $H_{1}$ is a genus-$g$ graph with $(\Delta+1)n$ vertices and maximum degree $2\Delta$. Observe also that $G$ is a uniform shallow minor of $H_{1}$ of depth $2$, and hence Lemma 3.9 shows that $\gamma^{(1)}(G)\lesssim(\Delta+1)\gamma^{1}(H_{1})$.
@@
For the remaining lemmas, we need only consider maximal graphs of genus $g$ (triangulations).
@@ 4 Bounds via $1$-dimensional Extremal $L_{p}$-Embedded Spread
Note that we can avoid the constraints on $n$ for the bounds in Section 2.2 since $\gamma^{(1)}(G)\leq 2+\frac{2}{n-1}\lesssim 1$: If $n$ is even, assign $y$ to be $\frac{2}{n}$ for every vertex, and assign $f$ to be $-1$ for half the vertices and $1$ for the other half of the vertices (this results in the stronger bound of $\gamma^{(1)}(G)\leq 2$). If $n$ is odd (and at least $3$), assign $y$ to be $\frac{2}{n-1}$ for every vertex, and assign $f$ to be an even split of $-1$ and $1$ for all but one vertex, which is assigned 00 (resulting in the stated bound).

### v2

@@ 1 Introduction
$K_{h}$-minor-free $\operatorname{\mathcal{O}}(h\sqrt{n})$ $\operatorname{\mathcal{O}}(n^{1+\varepsilon})$ [^2] Assumes $h$ is a constant, and uses a tower function in terms of $h$. Note that some more tractable algorithms are known with weaker guarantees [WN11]. [KR10]
@@ 1.1.2 Reweighted Spectral Partitioning
One useful bound on $\lambda_{2}^{*}(G)=\gamma^{(n)}\leq\gamma^{(1)}(G)$ is:
Observation 1.3.
For any graph $G$, $\gamma^{(1)}(G)\lesssim 1$.
Proof.
If $n$ is even, assign $y$ to be $\frac{2}{n}$ for every vertex, and assign $f$ to be $-1$ for half the vertices and $1$ for the other half of the vertices, so then $\gamma^{(1)}(G)\leq 2\lesssim 1$. If $n$ is odd (and at least $3$), assign $y$ to be $\frac{2}{n-1}$ for every vertex, and assign $f$ to be an even split of $-1$ and $1$ for all but one vertex, which is assigned 00, resulting in $\gamma^{(1)}(G)\leq 2+\frac{2}{n-1}\lesssim 1$. ∎
@@ 2.1 Bounds on $\lambda_{2}(G)$ for Special Graph Classes
For a triangulated genus-$g$ graph $G$ with $n$ vertices and maximum degree $\Delta$, the second smallest eigenvalue of the Laplacian matrix of $G$ has $\lambda_{2}(G)\lesssim\frac{\text{poly}(\Delta)\cdot g}{n}$, where $\text{poly}(\Delta)$ denotes a polynomial in $\Delta$.
@@ 3.1 Geometric Intersection Graphs and Nearest-Neighbour Graphs
The above two statements apply algorithmically to reweighted spectral partitioning. We note that, in some applications, the geometric information may actually be known, in which case the constructions above can be used more directly:
Corollary 3.8.
Let $G$ be a $d$-dimensional $k$-ply neighbourhood system with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and radii for each vertex. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
Corollary 3.9.
Let $G$ be a $d$-dimensional $k$-nearest neighbour graph with maximum degree $\Delta$, provided as a set of $d$-dimensional coordinates and a value $k$. Then, in polynomial time, we can compute a balanced $2/3$-vertex-separator of size $\operatorname{\mathcal{O}}\left(\sqrt{\min\{d,\log\Delta\}}\cdot\left(\frac{k}{n}\right)^{\frac{1}{d}}\right)$.
@@ 3.2 Bounded-Genus Graphs
Kelner’s bound on $\lambda_{2}$ (stated in Theorem 2.3) for triangulated genus-$g$ graphs also makes use of circle packings [Kel04, Kel06b, Kel06a], and, as we will see, can also be generalized to a bound on $\lambda_{2}^{*}$ and $\gamma^{(1)}$. However, the resulting bound also has a (large) polynomial dependence on the maximum degree $\Delta$. Rather than directly applying this bound, we will apply a new transformation that allows us to almost fully tighten this dependence. Specifically, our new transformation will take a genus-$g$ graph with $n$ vertices and maximum degree $\Delta$, and produce a graph with $O(\Delta\cdot n)$ vertices and constant maximum degree. We will then relate the values of $\gamma^{(1)}$ for the two graphs. In the case of $\lambda_{2}^{*}$ and $\gamma^{(1)}$, this relationship will result in a greatly improved $O(\log\Delta)$ factor for the upper bound.
@@
Recall that a graph $G$ has genus $g$ if it has a combinatorial embedding (represented as a rotation system) onto a surface of genus $g$. That is, each vertex has a circular ordering on its vertices, forming a locally-planar embedding. Such an embedding is cellular if each face induced by the rotation system is simple. A cellular embedding is a triangulation if each face defined by the rotation system is a triangle.
It is notable that Kelner’s bound on $\lambda_{2}$ does not apply to non-triangulated genus-$g$ graphs, despite the paper claiming otherwise. The claimed argument is as follows:
  • Adding edges to a graph can only increase $\lambda_{2}$.
  • Add edges until the graph is maximal.
  • Apply the bound on triangulated genus-$g$ graphs that Kelner proves in the rest of his paper. Note that this bound relies on a generalized circle packing theorem of He and Schramm [HS93] that applies only to simple triangulations of genus-$g$ surfaces.
This argument has two separate issues:
  • A maximal genus-$g$ graph is not always a triangulation. In particular, $K_{5}$ is a maximal genus-$1$ graph that is not triangulated in any embedding into the torus. We may prove this by contradiction: Suppose this was not the case, and that it had a triangulated embedding in the torus. Then, by the Euler characteristic, such an embedding must have $5$ faces. We may count the number of edge-face incidences. Since the embedding is triangulated (and hence cellular), each edge is incident to exactly two faces, and there are $10$ edges and hence $2\cdot 10=20$ incidences in total. Since the embedding is triangulated, there are exactly $3$ edges incident to each face, and hence there are $3\cdot 5=15$ incidences in total – a contradiction.
  • Naively adding edges may increase the maximum degree, possibly up to $\Omega(n)$, and the bound that Kelner obtains for triangulations has a dependence on the maximum degree.
It is worth noting that neither of these issues arises for planar graphs: Maximal planar graphs are always planar triangulations, and there is a known result that triangulates a planar graph while preserving the maximum degree [KB92, KB97]. For genus $g>0$, we will present a method to bypass these issues with carefully chosen Steiner vertices while tightening and generalizing the bound to $\lambda_{2}^{*}$ and $\gamma^{(1)}$.
Henceforth, we will fix a rotation system for our initial graph $G$. This rotation system need not be cellular, but we will assume that $G$ is simple (as we do throughout this paper). Importantly for our purposes, this will allow us to make certain local transformations to a graph without changing its genus, so long as the rotation system is respected in these transformations. These transformations will make use of the following structure:
@@
We start with a trivial case whose elimination we will use later: If $\gamma^{(1)}(H)\cdot r\cdot L\gtrsim 1$, then the result holds since $\gamma^{(1)}(G)\lesssim 1$ (see 1.3).
@@
We will be able to leverage this lemma in order to turn a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for bounded-genus graphs of constant maximum degree into a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for bounded-genus graphs of arbitrary maximum degree at only a $\log\Delta$ loss. In particular, we will obtain a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for triangulated bounded-genus graphs of constant degree by adapting Kelner’s techniques [Kel04, Kel06b, Kel06a] for bounding $\lambda_{2}$. Before exploring these techniques, we first show how to reduce to the case of triangulated bounded-genus graphs, which will use a similar technique:
@@
Let $G$ be a graph of genus $g$ with $n$ vertices and maximum degree $\Delta$. Then there is a triangulated genus-$g$ graph $H$ with $(\Delta+1)\cdot n$ vertices and maximum degree $O(\Delta)$, so that $\gamma^{(1)}(G)\lesssim(\Delta+1)\cdot\gamma^{(1)}(H)$.
@@
Let $H_{1}$ be the graph that adds $\Delta$ vertices around each vertex to $G$, and ensure that between any two edges in the rotation system around a vertex in $G$, at least one such edge is added. Observe that $H_{1}$ is a genus-$g$ graph with $(\Delta+1)n$ vertices and maximum degree $2\Delta$. Observe also that $G$ is a uniform shallow minor of $H_{1}$ of depth $2$, and hence Lemma 3.11 shows that $\gamma^{(1)}(G)\lesssim(\Delta+1)\gamma^{(1)}(H_{1})$.
@@
For the remaining lemmas, we need only consider triangulated genus-$g$ graphs.
@@ 4 Bounds via $1$-dimensional Extremal $L_{p}$-Embedded Spread
Note that we can avoid the constraints on $n$ for the bounds in Section 2.2 since $\gamma^{(1)}(G)\lesssim 1$ (see 1.3).